### v1

@@ 1 Introduction
$\overline{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}.$
@@
When $f$ is also differentiable, basic technical conditions (see e.g. [13, Theorem 11.39]) allow the primal and dual solution to be simultaneously obtained by solving a system of nonlinear equations characterizing the stationarity of the Lagrangian (10):
@@ 2.1 One-dimensional raking
$f(\beta;y)=\sum_{i=1}^{p}\beta_{i}\log\left(\frac{\beta_{i}}{y_{i}}\right)-\beta_{i}+y_{i},\qquad f^{*}(z)=y^{T}\left(e^{z}-1\right)\text{ and }\nabla_{z}f^{*}\left(z\right)=y\odot e^{z}.$
@@ 2.2 Two-dimensional raking and iterative proportional fitting (IPF) algorithm
(25) $\min_{\beta}\max_{\lambda_{1},\lambda_{2}}\mathcal{L}\left(\beta,\lambda_{1},\lambda_{2}\right):=f\left(\beta;y\right)+\lambda_{1}^{T}\left(\beta^{T}1-s_{1}\right)+\lambda_{2}^{T}\left(\beta 1-s_{2}\right),$
@@
The IPF algorithm is a particular way to exploit the primal-dual relationship (23)-(26) and correspondence (27) , but it breaks down when we change the optimization problem (23) by modifying the objective $f$, including adding weights, or incorporating additional terms. We use algorithms that attack the dual directly in more general formulations.
@@ 2.3 Multi-dimensional weighted raking
Here, we study more in detail the general case with $k$ auxiliary variables which are a linear combination of the data. We suppose that we have a vector of detailed observations $y\in\mathbb{R}^{p}$. The margin vector $\sum_{i\in\mathcal{U}}x_{i}$ is denoted $s$ and the constraints become $A\beta=s$. We introduce a vector of weights $q\in\mathbb{R}^{p}$ that do not depend on $y$ or $s$ and denote $\frac{1}{q}=\left(\frac{1}{q_{1}},\cdots,\frac{1}{q_{i}},\cdots,\frac{1}{q_{p}}\right)^{T}$ the vector containing the inverses of the weights. Denoting
@@
(31) $\beta^{*}=\operatorname*{arg\,min}_{\beta}f_{1/q}\left(\beta;y\right)\quad\mbox{s.t.}\quad A\beta=s$
@@
(33) $\min_{\lambda}\lambda^{T}s+\sum_{i}\frac{1}{q_{i}}f_{i}^{*}\left(-q_{i}a_{i}^{T}\lambda\right).$
Taking the derivative with respect to $\lambda$, we get the dual optimality conditions:
@@ 3 Uncertainty
A review on methods to use priors on the margins to compute the posterior distribution for the total $\mathcal{Y}$ is given in  [14]. However, in this paper, the total sum $\mathcal{Y}$ and its variance are known and we are interested in using the known variances and covariances on the observations $y$ and the margins $s$ to estimate the unknown covariance matrix of the raked values $\beta$.
@@
Theorem 3.1 (Covariance of the raked values $\beta^{*}$).
Let $\theta$ denote the expectation of the observations and margins, $\Sigma$ denote their covariance matrix, and $\phi^{\prime}_{\theta}$ denote the matrix of the partial derivatives of $\phi$ with respect to the observations $y$ and margins $s$ taken at $\theta$. Then the covariance matrix of the raked values $\beta^{*}$ is given by
$\Sigma_{\beta}=\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}.$
@@
Suppose that we have $n$ vectors of observations $\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$ taking values in $\mathbb{R}^{p+k}$. We can compute their sample mean $\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}$ and through the central limit theorem, we have
@@
where $\theta=\mathbb{E}\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$, $\Sigma$ is the covariance matrix
(61) $\Sigma=\begin{pmatrix}\Sigma_{y}&\Sigma_{ys}\\ \Sigma_{ys}&\Sigma_{s}\end{pmatrix}$
and $\mathcal{N}\left(0,\Sigma\right)=T$ denotes the multivariate normal distribution with mean 00 and covariance matrix $\Sigma$. If the function $\phi$ mapping the detailed observations and the aggregate observations to the raked estimates is differentiable at $\theta$, then we have
(62) $\sqrt{n}\left(\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{n}\end{matrix}\right)-\phi\left(\theta\right)\right)\rightarrow\phi^{\prime}_{\theta}\left(T\right)$
@@
(63) $\phi^{\prime}_{\theta}\left(T\right)=\begin{pmatrix}\frac{\partial\phi_{1}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{1}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial s_{k}}\left(\theta\right)\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial\phi_{p}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial s_{k}}\left(\theta\right)\\ \frac{\partial\phi_{p+1}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+1}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p+1}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+1}}{\partial s_{k}}\left(\theta\right)\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial\phi_{p+k}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+k}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p+k}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p+k}}{\partial s_{k}}\left(\theta\right)\end{pmatrix}T,$
and so $\sqrt{n}\left(\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{n}\end{matrix}\right)-\phi\left(\theta\right)\right)$ converges weakly to a multivariate normal distribution with mean 00 and covariance $\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}$. To relate the general result to the specific example (56), with the $\chi^{2}$ distance, (57) is
(64) $\begin{pmatrix}I&\left[\text{diag}\left(q\odot y\right)\right]A^{T}\\ A&0\end{pmatrix}\begin{pmatrix}\beta\\ \lambda\end{pmatrix}=\begin{pmatrix}y\\ s\end{pmatrix}.$
If $\beta_{1}^{*}$ solves $F\left(\beta,\lambda;y_{1},s_{1}\right)=0$ and $\beta_{2}^{*}$ solves $F\left(\beta,\lambda;y_{2},s_{2}\right)=0$, then $\beta_{1}^{*}+\alpha\beta_{2}^{*}$ solves
$F\left(\beta,\lambda;y_{1}+y_{2},s_{1}+\alpha s_{2}\right)=0.$
From this, we can conclude that
(65) $\phi\left(\begin{matrix}\bar{y}_{n}\\ \bar{s}_{s}\end{matrix}\right)=\frac{1}{n}\sum_{i=0}^{n}\phi\left(\begin{matrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{matrix}\right).$
From  (59), the covariance matrix of $\phi\left(\begin{matrix}y^{(i)}\\ s^{(i)}\end{matrix}\right)$ is equal to the covariance matrix of the raked values $\Sigma_{\beta}$. Thus we have through the central limit theorem
(66) $\sqrt{n}\left(\frac{1}{n}\sum_{i=0}^{n}\phi\left(\begin{matrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{matrix}\right)-\phi\left(\theta\right)\right)\rightarrow\mathcal{N}\left(0,\Sigma_{\beta}\right).$
Gathering  (62) and  (66), we finally get
(67) $\Sigma_{\beta}=\phi^{{}^{\prime}T}_{\theta}\Sigma\phi^{\prime}_{\theta}.$
@@
Due to the choice of the distance function $f$, the Hessian matrices $\nabla^{2}_{\beta}f\left(\beta;y\right)$ and $\nabla^{2}_{\beta y}f\left(\beta;y\right)$ are diagonal. If we prune the constraints to ensure $A$ has full rank as described above, both $D_{\beta,\lambda}F$ and $D_{y,s}F$ are invertible. The system  (72) has then a unique solution.
@@ 4 Simulated Example
We apply the raking procedure to a small synthetic 2-dimensional example with two categorical variables $X_{1}$ and $X_{2}$. We suppose that $X_{1}$ can take $m=3$ values $\left[1;2;3\right]$ and that $X_{2}$ can take $n=5$ values $\left[1;2;3;4;5\right]$ and we generate $m\times n$ values following a uniform distribution over the interval $\left[2;3\right]$ to fill the initial matrix $\beta_{0}$. We compute the margins $s_{1}=\beta_{0}^{T}1$ and $s_{2}=\beta_{0}1$. Then we add some random noise to get the observations $y_{0}=\beta_{0}+\mathcal{N}\left(\mu=0,\sigma=0.1\right)$. We then generate $N=1000$ samples following a multivariate normal distribution with mean $y_{0}$ and covariance $\Sigma$. The off-diagonal elements of $\Sigma$ are equal to $0.01$ while the diagonal elements of $\Sigma$ are equal to $\Sigma_{k,k}=0.1\times k$ for $k=1,\cdots,m\times n$. In this particular example, we suppose that there is no uncertainty on the margins. We can then compute the mean $\bar{y}$ of the observations $y^{(i)}$ and apply the raking procedure to the mean to get $\beta^{*}=\phi\left(\bar{y},s_{1},s_{2}\right)$. The results are shown in Figure 1. The blue dots represent the initial observations while the orange squares represent the raked values. The uncertainty on the observations and the raked values is marked by a segment of length two times the standard deviation and centered on the observations or the raked values.
Figure 1: Initial observations (blue dots) and corresponding raked values (orange squares) with their corresponding uncertainties. The raked values are close to the initial observations, but have lower uncertainties as there is no uncertainty on the margins.
@@
Finally, we compare in Figure 4 the raked value $\beta^{*}$ obtained when raking the mean of the $1000$ samples and when raking each of the samples and computing the mean of the raked values. We also compare the variance of $\beta^{*}$ computed with the uncertainty propagation method described in section 3 with the variance computed by raking the $1000$ samples of the observations and computing the sample variance of the corresponding raked values (that is, using an empirical process to quantify uncertainty). We get very similar results using both methods, showing that the asymptotic method leveraging the implicit function theorem is very effective.
Figure 4: Mean (left) and variance (right) of the raked values computed with the delta method or with all the samples. The method we developed gives nearly identical results as the Monte Carlo method.
@@ 5 Application
Our dataset [9] contains for each US state, year, sex, and age group the estimated number of deaths by cause, race and ethnicity and county. These estimates are raked to aggregate estimates, typically all-race state-wide cause-specific and all-cause counts, while trying to preserve key relational information in the detailed estimates, as described in detail below.
There are $I=3$ different causes (communicable, maternal, neonatal, and nutritional diseases (Comm.); non-communicable diseases (NCD); and injuries (Inj.)) and $J=5$ different races (White; Black; American Indian or Alaska Native (AIAN); Asian or Pacific Islander (API); and Hispanic/Latino (Hisp.)) in the dataset. We denote $y_{i,j,k}$ the number of deaths for cause $i=1,\cdots,I$, race $j=1,\cdots,J$ and county $k=1,\cdots,K$. We denote $i=0$ for the sum of deaths over all causes, $j=0$ for the sum of deaths over all races, and $k=0$ for the sum of deaths over all counties, that is the number of deaths at the state level. The only aggregate data known are provided by the GBD  [2] and are the total number of deaths for the state $s_{0}=y_{0,0,0}$ and the total number of deaths due to cause $i$ for the state $s_{i}=y_{i,0,0}$.
The modeling challenge here is to also ensure internal consistency, guaranteeing that:
@@
  Minimize $\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)$ under the constraints $\sum_{k=1}^{K}\beta_{i,0,k}=s_{i}$ for all $i=1,\cdots,I$ and $\sum_{i=1}^{I}\beta_{i,0,k}=\beta_{0,0,k}$ for all $k=1,\cdots,K$.
@@
     Minimize $\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$ under the constraints $\sum_{i=1}^{I}\beta_{i,j,k}=\beta_{0,j,k}$ for all $j=1,\cdots,J$ and $\sum_{j=1}^{J}\beta_{i,j,k}=\beta_{i,0,k}$ for all $i=1,\cdots,I$.
@@
Using the formulations and methods in the paper, we can rake on cause, race and county in one step, using any distance function we like, using weighting as necessary, and obtain uncertainty through a single solve, avoiding re-solving the problem 1000 times (once for each draw). The corresponding workflow is described in Algorithm 2.
@@
  Minimize $\sum_{k=1}^{K}f\left(\beta_{0,0,k},y_{0,0,k}\right)+\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)+\sum_{k=1}^{K}\sum_{j=1}^{J}f\left(\beta_{0,j,k},y_{0,j,k}\right)+\sum_{k=1}^{K}\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$ under the constraints:
@@
In this proposed workflow, if we denote $l=(I+1)(J+1)(k-1)+(I+1)j+i+1$ and we take $s_{l}=0$ for $l>i$, we can rewrite the problem as
@@
where $p=(I+1)(J+1)K$ and $A$ is a $k\times p$ sparse matrix containing only 00s, $1$s and $-1$s. In the initial formulation of Algorithm 2, there are $1+I+\left(I+J+2\right)K$ constraints. However, the first constraint $\sum_{k=1}^{K}\beta_{0,0,k}=s_{0}$ and the last $K$ constraints $\sum_{j=1}^{J}\beta_{I,j,k}-\beta_{I,0,k}=0$ are a linear combination of the other constraints. We can drop them and finally get a $k\times p$ matrix $A$ that has full rank, with $k=I+\left(I+J+1\right)K$. All the linear systems in Sections 2 and 3 will then have a unique solution.
In Figure 5, we compute the raking procedure for Delaware, both sexes, and age group 25-to-30-year-old, using the entropic distance and equal weights. We compare the reported values of the death rates (number of deaths divided by the population for each race and county) with the corresponding raked values for both workflows for all 72 cases (3 counties, 5 races and all races, 3 causes and all causes). The 1-step approach gives raked values slightly closer to the initial values than the 4-steps workflow, meaning that the raked estimates stay closer to the original reported values while satisfying the requirements. This improvement is easier to see for larger reported death rates on the linear scale plot in Figure 5.
Figure 5: Initial rate of death and corresponding raked values for the two proposed workflows. The raked values obtained with the proposed 1-step workflow are usually closer to the initial observations than the raked values obtained with the previous 4-step workflow.
We now compare the raked values with their associated uncertainty to the initial values. To better compare the different causes, races and counties on the same scale, we compare mortality rates. In Figure 6, Figure 7 and Figure 8, we look at the initial values and the raked values for Delaware, both sexes and age group 0-to-1-year-old. The uncertainty is represented by a segment of length two standard deviations and centered on the estimated value. Races $3$ and $4$ with much smaller population numbers have the largest uncertainties for both the initial and the raked values. This uncertainty on the raked values is significantly bigger than the uncertainty on the initial values. The initial uncertainty on the margins $s$ (that is on the GBD values) are much larger than the uncertainties on the observations $y$, and as a result the raking process significantly increases the final uncertainty on these raked values.

### v2

@@ 1 Introduction
$\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}.$
@@
When $f$ is also differentiable, basic technical conditions (see e.g. [16, Theorem 11.39]) allow the primal and dual solutions to be simultaneously obtained by solving a system of nonlinear equations characterizing the stationarity of the Lagrangian (10):
@@ 2.1 One-dimensional raking
$f(\beta;y)=\sum_{i=1}^{p}\beta_{i}\log\left(\frac{\beta_{i}}{y_{i}}\right)-\beta_{i}+y_{i},\qquad f^{*}(z)=y^{T}\left(e^{z}-1\right)\text{ and }\nabla_{z}f^{*}\left(z\right)=y\odot e^{z},$
@@ 2.2 Two-dimensional raking and iterative proportional fitting (IPF) algorithm
(25) $\mathcal{L}\left(\beta,\lambda_{1},\lambda_{2}\right):=f\left(\beta;y\right)+\lambda_{1}^{T}\left(\beta^{T}1-s_{1}\right)+\lambda_{2}^{T}\left(\beta 1-s_{2}\right),$
@@
The IPF algorithm is a particular way to exploit the primal-dual relationship (23)-(26) and correspondence (27) , but the classic form breaks down when we change the optimization problem (23) by modifying the objective $f$, including relative weights, or incorporating additional terms. In all of these formulations, we found it far easier to solve the dual directly, and then use the dual solution to recover the primal at the end.
@@ 2.3 Multi-dimensional weighted raking
Here, we study more in detail the general case with $k$ auxiliary variables which are a linear combination of the data. We suppose that we have a vector of detailed observations $y\in\mathbb{R}^{p}$. The margin vector $\sum_{i\in\mathcal{U}}x_{i}$ is denoted $s$ and the constraints become $A\beta=s$. For example, in the case of 2D raking described above, we can use classic kronecker identities to rewrite the row and column sum constraints as a single linear system on the vectorized $\beta$:
$1\beta$ $=s_{1}^{T}\quad\Rightarrow\quad I\otimes 1\mbox{vec}(\beta)=s_{1}$
$\beta 1$ $=s_{2}\quad\Rightarrow\quad 1^{T}\otimes I\mbox{vec}(\beta)=s_{2}$
$A$ $=\begin{bmatrix}I\otimes 1\\ 1^{T}\otimes I\end{bmatrix},\quad s=\begin{bmatrix}s_{1}\\ s_{2}.\end{bmatrix}$
This representation is less compact than (23), which explicitly uses the matrix structure of the decision variable $\beta$ in the 2D case.
Next, we introduce a vector of weights $q\in\mathbb{R}^{p}$ that do not depend on $y$ or $s$ and denote $\frac{1}{q}=\left(\frac{1}{q_{1}},\cdots,\frac{1}{q_{i}},\cdots,\frac{1}{q_{p}}\right)^{T}$ the vector containing the inverses of the weights. Denoting
@@
(31) $\min_{\beta}f_{1/q}\left(\beta;y\right)\quad\mbox{s.t.}\quad A\beta=s$
@@
(33) $\min_{\lambda}\lambda^{T}s+\sum_{i}\frac{1}{q_{i}}f_{i}^{*}\left(-q_{i}a_{i}^{T}\lambda\right)$
where $a_{i}$ denotes the $i$th column of $A$. Taking the derivative with respect to $\lambda$, we get the dual optimality conditions:
@@ 3 Uncertainty
[13] used the uncertainty on the raking weights to compute the variance of the resulting estimate.  [12] accounted for the raking procedure to estimate the variance of the estimator.  [14] used bootsrap method to estimate the uncertainty of the raking estimator for small area estimation. A review on methods to use priors on the margins to compute the posterior distribution for the total $\mathcal{Y}$ is given in  [17]. However, in this paper, the total sum $\mathcal{Y}$ and its variance are known and we are interested in using the known variances and covariances on the observations $y$ and the margins $s$ to estimate the unknown covariance matrix of the raked values $\beta$.
@@
Theorem 3.1 (Covariance of the raked values).
Suppose that the vector of observations and margins $\begin{pmatrix}y\\ s\end{pmatrix}$ is a random vector with expectation $\theta$ and covariance matrix $\Sigma$. Given $n$ samples of $\begin{pmatrix}y\\ s\end{pmatrix}$, denoted $\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$ for $i=1,\cdots,n$, denote their sample mean by $\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}$. Let $\phi^{\prime}_{\theta}$ denote the matrix of the partial derivatives of $\phi$ with respect to the observations $y$ and margins $s$ taken at $\theta$. If $\beta^{*}_{n}$ denotes the raked values of the sample mean, then the covariance matrix of the raked values $\beta^{*}_{n}$ is given by
$\Sigma_{\beta_{n}}=\phi^{\prime}_{\theta}\Sigma\phi^{{}^{\prime}T}_{\theta}.$
@@
Given $n$ vectors of observations $\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}\in\mathbb{R}^{p+k}$ and their sample mean $\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}$, from the central limit theorem, we have
@@
where $\theta=\mathbb{E}\begin{pmatrix}y^{(i)}\\ s^{(i)}\end{pmatrix}$, and $\Sigma$ is the covariance matrix
(61) $\Sigma=\begin{pmatrix}\Sigma_{y}&\Sigma_{ys}\\ \Sigma_{ys}&\Sigma_{s}\end{pmatrix}.$
If the mapping $\phi:\mathbb{R}^{p+k}\rightarrow\mathbb{R}^{p}$ from detailed and aggregate observations to the raked estimates is differentiable at $\theta$, then we have
(62) $\sqrt{n}\left(\phi\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}-\phi\left(\theta\right)\right)\rightarrow\phi^{\prime}_{\theta}\left(T\right)$
@@
(63) $\phi^{\prime}_{\theta}\left(T\right)=\begin{pmatrix}\frac{\partial\phi_{1}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{1}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{1}}{\partial s_{k}}\left(\theta\right)\\ \vdots&&\vdots&\vdots&&\vdots\\ \frac{\partial\phi_{p}}{\partial y_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial y_{p}}\left(\theta\right)&\frac{\partial\phi_{p}}{\partial s_{1}}\left(\theta\right)&\ldots&\frac{\partial\phi_{p}}{\partial s_{k}}\left(\theta\right)\end{pmatrix}T,$
and so $\sqrt{n}\left(\phi\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix}-\phi\left(\theta\right)\right)$ converges weakly to a multivariate normal distribution with mean 00 and covariance $\phi^{\prime}_{\theta}\Sigma\phi^{{}^{\prime}T}_{\theta}$.
Theorem 3.1 extends the delta method to the higher dimensional setting, where in addition to the classic $p$ scoring equations we have $k$ additional equations with randomly distributed right hand sides to inform the estimate. In practice, social scientists often use draws, or samples of original estimates and the margins, to estimate the variance of the raked estimates. In this scheme, for each draw of observations and margins, the raking problem is solved, and then the mean and variance are computed from the output draws. To understand this process better, consider the raked values of the random vector $\begin{pmatrix}y\\ s\end{pmatrix}$. This is also a random vector, denoted $\phi\begin{pmatrix}y\\ s\end{pmatrix}$ with expectation $\mathbb{E}\phi\begin{pmatrix}y\\ s\end{pmatrix}$ and covariance matrix $\Sigma_{\beta}$. Given $n$ samples of the raked vector $\phi\begin{pmatrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{pmatrix}$, from the central limit theorem we have
(64) $\sqrt{n}\left(\frac{1}{n}\sum_{i=0}^{n}\phi\begin{pmatrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{pmatrix}-\mathbb{E}\phi\begin{pmatrix}y\\ s\end{pmatrix}\right)\rightarrow\mathcal{N}\left(0,\Sigma_{\beta}\right).$
If we have to first order that
(65) $\frac{1}{n}\sum_{i=0}^{n}\phi\begin{pmatrix}y^{\left(i\right)}\\ s^{\left(i\right)}\end{pmatrix}\approx\phi\begin{pmatrix}\bar{y}_{n}\\ \bar{s}_{n}\end{pmatrix},$
then (62) and  (64), together tells us that the draws process has the asymptotically correct statistics:
(66) $\mathbb{E}\phi\begin{pmatrix}y\\ s\end{pmatrix}\approx\phi\left(\theta\right)\text{ and }\Sigma_{\beta}\approx\phi^{\prime}_{\theta}\Sigma\phi^{{}^{\prime}T}_{\theta}.$
The approximation (65) trivially holds when $\phi$ is a linear map, for example, the optimality conditions in a linear regression problem. However, whenever $\phi$ is nonlinear, or indeed even under affine constraints, the conclusion may not hold. For the raking problem, we show empirically in the experiments that given enough draws we actually can get close to the asymptotic statistics, but we incur a systematic bias when we use the draws.
3.1 Computing Derivatives
@@
Due to the choice of the distance function $f$, the Hessian matrices $\nabla^{2}_{\beta}f\left(\beta;y\right)$ and $\nabla^{2}_{\beta y}f\left(\beta;y\right)$ are diagonal. If we prune the constraints to ensure $A$ has full rank, both $D_{\beta,\lambda}F$ and $D_{y,s}F$ are invertible. The system  (71) then has a unique solution.
@@ 4 Simulated Example
We apply the raking procedure to a small synthetic 2-dimensional example with two categorical variables $X_{1}$ and $X_{2}$. We suppose that $X_{1}$ can take $m=3$ values $\left[1;2;3\right]$ and that $X_{2}$ can take $n=5$ values $\left[1;2;3;4;5\right]$ and we generate $m\times n$ values following a uniform distribution over the interval $\left[2;3\right]$ to fill the initial matrix $\beta_{0}$. We compute the margins $s_{1}=\beta_{0}^{T}1$ and $s_{2}=\beta_{0}1$. We add random noise to get the observations $y_{0}=\beta_{0}+\mathcal{N}\left(\mu=0,\sigma=0.1\right)$. We generate $N$ samples following a multivariate normal distribution with mean $y_{0}$ and covariance $\Sigma$. The off-diagonal elements of $\Sigma$ are equal to $0.01$ while the diagonal elements of $\Sigma$ are equal to $\Sigma_{k,k}=0.1\times k$ for $k=1,\cdots,m\times n$. In this particular example, we suppose that there is no uncertainty on the margins. We can then compute the mean $\bar{y}$ of the observations $y^{(i)}$ and apply the raking procedure to the mean to get $\beta^{*}=\phi\left(\bar{y},s_{1},s_{2}\right)$. The results are shown in Figure 1. The blue dots represent the initial observations while the orange squares represent the raked values. The uncertainty on the observations and the raked values is marked by a segment of length two times the standard deviation and centered on the observations or the raked values.
Figure 1: Initial observations (blue dots) and corresponding raked values (orange squares) with their corresponding uncertainties. The raked values are close to the initial observations, but have lower uncertainties as there is no uncertainty on the margins, and marginal constraints help to inform the estimates.
@@
Finally, we compare in Figure 4 and Figure 5 the raked value $\beta^{*}$ obtained when raking the mean of the $N$ samples and when raking each of the samples and computing the mean of the raked values, for $N=1000,10000,100000$, shown in Figure 4. The means of the raked draws are close to the raked mean, even for a small number of samples.
We also compare the variance of $\beta^{*}$ computed with the uncertainty propagation method described in section 3 with the variance computed by raking the samples of the observations for $N=1000,10000,100000$, and computing the sample variance of the corresponding raked values (that is, using an empirical process to quantify uncertainty). The comparison is shown in Figure 5. We see more variability in the alignment of the variance estimates for low samples, and a very slight bias in the Monte Carlo process at high numbers of samples compared to the delta method.
Figure 4: Mean of the raked values computed with the new delta method or with all the samples for $N=1000,10000,100000$ (left to right). The new delta method gives nearly identical results as the Monte Carlo method, which is accurate even for 1000 samples.
Figure 5: Variance of the raked values computed with the delta method or with all the samples for $N=1000,10000,100000$ (left to right). The new delta method is close to the Monte Carlo method when the latter uses many draws. More variation is evident in the Monte Carlo results for fewer samples.
@@ 5 Application
Our dataset [9] contains the estimated number of deaths by cause, race/ ethnicity and county for each US state, year, sex, and age group. These estimates are raked to aggregate estimates, typically all-race state-wide cause-specific and all-cause counts, while trying to preserve key relational information in the detailed estimates, as described below.
In the dataset, there are $I=3$ different causes: communicable, maternal, neonatal, and nutritional diseases (Comm.), non-communicable diseases (NCD), and injuries (Inj.) and $J=5$ different race/ethnicity groups: White, Black, American Indian or Alaska Native (AIAN), Asian or Pacific Islander (API), and Hispanic/Latino (Hisp.). We denote $y_{i,j,k}$ the number of deaths for cause $i=1,\cdots,I$, race $j=1,\cdots,J$ and county $k=1,\cdots,K$. We use $i=0$ for the sum of deaths over all causes, $j=0$ for the sum of deaths over all races, and $k=0$ for the sum of deaths over all counties, i.e. the number of deaths at the state level. Aggregate data are provided by the GBD  [2] and are the total number of deaths for the state $s_{0}=y_{0,0,0}$ and the total number of deaths due to cause $i$ for the state $s_{i}=y_{i,0,0}$.
The modeling challenge in the application is to ensure internal consistency:
@@
  Minimize $\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)$ under the constraints
$\sum_{k=1}^{K}\beta_{i,0,k}=s_{i}\;\forall i=1,\cdots,I,\quad\sum_{i=1}^{I}\beta_{i,0,k}=\beta_{0,0,k}\;\forall k=1,\cdots,K.$
@@
     Minimize $\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$ under the constraints
$\sum_{i=1}^{I}\beta_{i,j,k}=\beta_{0,j,k}\;\forall j=1,\cdots,J,\quad\sum_{j=1}^{J}\beta_{i,j,k}=\beta_{i,0,k}\;\forall i=1,\cdots,I.$
@@
Using the formulations and methods in the paper, we can rake on cause, race and county in one step, with a choice of distance function, weighting as necessary, and get uncertainty estimates using a single solve, rather than re-solving for each draw. The corresponding workflow is described in Algorithm 2.
@@
  Minimize
$\sum_{k=1}^{K}f\left(\beta_{0,0,k},y_{0,0,k}\right)$ $+\sum_{i=1}^{I}\sum_{k=1}^{K}f\left(\beta_{i,0,k},y_{i,0,k}\right)+\sum_{k=1}^{K}\sum_{j=1}^{J}f\left(\beta_{0,j,k},y_{0,j,k}\right)$
$+\sum_{k=1}^{K}\sum_{i=1}^{I}\sum_{j=1}^{J}f\left(\beta_{i,j,k},y_{i,j,k}\right)$
under the constraints:
@@
Denoting $l=(I+1)(J+1)(k-1)+(I+1)j+i+1$ and taking $s_{l}=0$ for $l>i$, we can rewrite the problem as
@@
where $p=(I+1)(J+1)K$ and $A$ is a $k\times p$ sparse matrix containing with entries in $\{-1,0,1\}$. In the initial formulation of Algorithm 2, there are $1+I+\left(I+J+2\right)K$ constraints. However, the first constraint $\sum_{k=1}^{K}\beta_{0,0,k}=s_{0}$ and the last $K$ constraints $\sum_{j=1}^{J}\beta_{I,j,k}-\beta_{I,0,k}=0$ are a linear combination of the other constraints. We can drop them and finally get a $k\times p$ matrix $A$ that has full rank, with $k=I+\left(I+J+1\right)K$. All the linear systems in Sections 2 and 3 then have a unique solution.
In Figure 6, we compute the raking procedure for Delaware, both-sex, age group 25-to-30 yearolds, using entropic distance and equal weights. We compare the reported values of the death rates (number of deaths divided by the population for each race and county) with the corresponding raked values for both workflows for all 72 cases (3 counties, 5 races and all races, 3 causes and all causes). The 1-step approach gives raked values slightly closer to the initial values than the 4-steps workflow, meaning that the raked estimates stay closer to the original reported values while satisfying the requirements. This improvement is easier to see for larger reported death rates on the linear scale plot in Figure 6.
Figure 6: Initial rate of death and corresponding raked values for the two proposed workflows for Delaware, age group 25-30. The raked values obtained with the proposed 1-step workflow are usually closer to the initial observations than the raked values obtained with the previous 4-step workflow.
We now compare the raked values with their associated uncertainty to the initial values. To better compare the different causes, races and counties on the same scale, we compare mortality rates. In Figure 7, Figure 8 and Figure 9, we look at the initial values and the raked values for both-sex Delaware, age group 0-to-1 yearolds. The uncertainty is represented by a segment of length two standard deviations and centered on the estimated value. Races $3$ and $4$ with much smaller population numbers have the largest uncertainties for both the initial and the raked values. This uncertainty on the raked values is significantly bigger than the uncertainty on the initial values. The initial uncertainty on the margins $s$ (that is on the GBD values) are much larger than the uncertainties on the observations $y$, and as a result the raking process significantly increases the final uncertainty on these raked values.